We had no power at our Haleakala site for two weeks, but last Friday, we got a portable generator up to the summit and got it hooked up. Our site manager spent the weekend testing all of the systems (enclosure, telescopes, instruments, etc.), and last night, we made a set of test observations with the 2m telescope. Science observations with the Spectral imager on the 2m telescope will resume tonight! We expect to make other telescopes/instruments available for science observations during the week.
At our Australia site, the 2m telescope was taken out-of-service on Feb 14 because of a malfunctioning amplifier card on one of the telescope's motors. That card was replaced over the weekend, and science observations resumed on the night of Feb 24. The FLOYDS spectrograph is still not available. The problem with FLOYDS was isolated to the camera. A replacement camera was installed last week, and tests are in progress.